# AlignMMBench: Evaluating Chinese Multimodal Alignment in Large Vision-Language Models

### v2

@@ Abstract
Evaluating the alignment capabilities of large Vision-Language Models (VLMs) is essential for determining their effectiveness as helpful assistants. However, existing benchmarks primarily focus on basic abilities using nonverbal methods, such as yes-no and multiple-choice questions. In this paper, we address this gap by introducing AlignMMBench, a comprehensive alignment benchmark specifically designed for emerging Chinese VLMs. This benchmark is meticulously curated from real-world scenarios and Chinese Internet sources, encompassing thirteen specific tasks across three categories, and includes both single-turn and multi-turn dialogue scenarios. Incorporating a prompt rewrite strategy, AlignMMBench encompasses 1,054 images and 4,978 question-answer pairs. To facilitate the evaluation pipeline, we propose CritiqueVLM, a rule-calibrated evaluator that exceeds GPT-4’s evaluation ability. Finally, we report the performance of representative VLMs on AlignMMBench, offering insights into the capabilities and limitations of different VLM architectures. All evaluation codes and data are available on https://alignmmbench.github.io/.
@@ 1 Introduction
Equipped with the Large Language Models (LLMs), Vision-Language Models (VLMs) demonstrate impressive performance in various visual tasks, such as image description and visual question answering [2, 35, 4]. Following alignment training, like Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), VLMs can comprehend and generate human language within visual contexts.
Recently, researchers in the Chinese community have made significant advancements and proposed multiple impressive VLMs, such as QwenVL [4], CogVLM [35], MiniCPM [16], Intern-VL [10], Intern-XComposer2V [11], Deepseek-VL [24], and Yi-VL [3]. To facilitate objective comparisons and evaluations among these models, researchers mainly utilize the following benchmarks: 1) general evaluation sets such as MME [13] and MMBench [22], and 2) domain-specific evaluation sets such as MathVista [25] and MMMU [41]. However, these benchmarks predominantly assess model capabilities through non-verbal methods and lack detailed evaluations of alignment performance, particularly in Chinese contexts. These limitations hinder a comprehensive understanding and optimization of emerging Chinese VLMs.
Table 1: Comparisons between AlignMMBench and other benchmarks.
@@
Ai2D [15] 4,903 English multi-choice Accuracy
LLaVABench [21] 150 English GPT Score
MathVista [25] 6,141 English [^1] MathVista contains 6.57% non-English questions, such as Chinese and Persian. GPT Accuracy
MME [13] 2,800 English yes/no Accuracy
MMBench [22] 3,217 Eng. & Chi. multi-choice Accuracy
MMMU [41] 11,500 English multi-choice Accuracy
MMStar [8] 1,500 English multi-choice Accuracy
MM-Vet [40] 205 English GPT Score
OCRBench [23] 1,000 English string match Accuracy
TouchStone [5] 908 English GPT Score
VisIT-Bench [6] 592 English GPT Elo / Win
AlignMMBench 4,978 Chinese CritiqueVLM Score
@@
Constructing a high-quality Chinese multimodal alignment benchmark is fully challenging. First, unlike the English context, which benefits from traditional benchmarks such as VQAv2 [14], TextVQA [33], and ChartQA [27], clean and publicly available Chinese multimodal corpora are exceedingly scarce. Therefore, it is necessary to collect and construct these resources from scratch. Second, the ambiguity inherent in the Chinese language is greater, requiring multiple annotators to repeatedly verify one single high-quality Chinese multimodal corpus. Nevertheless, building a Chinese multimodal benchmark is essential and promising. The characteristics of images and the underlying world knowledge can vary significantly between different languages. Consequently, relying solely on English datasets makes it difficult to comprehensively evaluate Chinese VLMs.
@@
To address this gap, we introduce AlignMMBench, a comprehensive evaluation benchmark specifically designed to assess the alignment capabilities of Chinese VLMs. AlignMMBench is meticulously curated from real-world scenarios and Chinese internet resources, encompassing thirteen specific tasks across three high-level categories, including both single-turn and multi-turn dialogue scenarios. Considering the diversity of query forms in real-world user scenarios, we introduce a LLM-based synonym expansion strategy, which transforms a single query into multiple stylistically different but semantically equivalent questions. As illustrated in Figure 1, AlignMMBench ultimately comprises 1,054 images and 4,978 question-answer pairs.
@@
Figure 1: Categories and examples of AlignMMBench. The chart on the left displays the categories of AlignMMBench, encompassing three main categories and thirteen specific tasks. The numbers below each category or task indicate the total number of images included. The right side of the pie chart presents two examples, illustrating instances from the incoherence and coherence categories.
@@
As the absence of fixed standard answers, the evaluation of open-ended questions is a significant and challenging research topic. One simple approach is to utilize the strongest LLM, such as GPT-4 [2], to score the model responses. However, the API-based GPT-4 operates as a closed-source black box, resulting in limited control for developers. Consequently, when the underlying model of the API undergoes changes, the corresponding scoring results are also affected. Considering these limitations, we fine-tuned an evaluator based on the open-source ChatGLM3-6B, named CritiqueVLM. Help with the detailed prompt rules and chain-of-thought reasoning, CritiqueVLM achieves higher consistency with human scoring compared to GPT-4.
@@
Based on AlignMMBench and CritiqueVLM , we evaluate ten popular VLMs that support Chinese. Our findings are as follows: 1) Most of VLMs can easily follow task instructions; 2) These VLMs excel in perception and understanding but perform poorly in reasoning and analysis; 3) Within dialogue contexts, VLMs exhibit suboptimal performance in the incoherence task compared to the coherence task, indicating that VLMs struggle to detect previous errors.
In summary, the main contributions of this paper are as follows:
@@
  • We propose AlignMMBench , a multimodal alignment benchmark that encompasses both single-turn and multi-turn dialogue scenarios. It includes three categories and thirteen capability tasks, with a total of 4,978 question-answer pairs. As far as we know, this is the public first alignment benchmark specifically designed for Chinese VLMs.
@@
  • To improve the controllability of evaluation scores, we introduce the CritiqueVLM , a ChatGLM3-6B based evaluator that has been rule-calibrated and carefully finetuned. With human judgements, its evaluation consistency surpasses that of GPT-4.
@@
  • We benchmark representative VLMs on AlignMMBench . Beyond their performance on AlignMMBench , we provide in-depth insights into the current state of Chinese VLM and highlight the areas that require further enhancement.
@@ 2 Related work
Early multimodal benchmarks predominantly focus on specific cross-modal tasks such as Image Caption [20, 30], Visual Grounding [18, 37], Visual Question Answering [32, 26, 28] and Optical Character Recognition [34]. The emergence of general VLMs underscores the necessity for comprehensive benchmarks capable of evaluating integrated capabilities.
To address these needs, recent works strive to propose various benchmarks containing complicated questions that expect VLMs to generate accurate and high-quality responses. These benchmarks can be categorized into two types based on their questions, including domain-specific and general scenarios. 1) Domain-specific benchmarks concentrate on particular subfields. For example, discipline reasoning [41, 25], Optical Character Recognition (OCR)  [23], chart understanding  [27, 15], and the hallucination test [21]. 2) General benchmarks provide multi-dimensional and diverse assessments, encompassing MME [13], MMBench [22], MMStar [8], MMVet [40], LLaVABench [8], VisIT-Bench [6], and TouchStone [5]. From another perspective, based on differences in result evaluation methods, the aforementioned benchmarks can be divided into deterministic and open-ended types. Deterministic benchmarks require the model to produce responses in a specific format through rule-based prompts. Common formats include multiple choice, yes / no, and strict matching. In contrast, open-ended benchmarks allow the model to generate unconstrained responses, which are then evaluated using advanced language models. Table 1 provides an overview of the categories and salient features of prominent benchmarks.
@@
In the realm of automated model response evaluation, employing a robust LLM as an evaluation assistant is widely adopted across various VLM benchmarks. MathVista [25] utilizes an LLM to extract answers from responses for more precise comparisons with reference answers, while other works [21, 40, 5, 6] adopt the LLM-as-a-Judge [43] approach. This method prompts the LLM to act as a judge, scoring responses or selecting the better response between two candidates. Existing VLM benchmarks use LLM-as-a-Judge methods with a unified evaluation prompt across all tasks and leverage GPTs [2] as their evaluator.
However, evaluating numerous models within a large benchmark can be costly and inefficient, with results potentially becoming unstable due to API-based model updates. In this work, we design a set of task-level evaluation prompts to improve the accuracy of LLM scoring. Additionally, we fine-tune ChatGLM3-6B [12] using a high-quality, human-annotated grading corpus to enhance the evaluation capabilities of our model.
@@ 3.1 Dataset composition
Figure 2: Overall framework of our work. AlignMMBench: Images are collected from real-world requests and Chinese websites. Following manual taxonomy and LLM-based question expansion, annotators generate reference answers within task-specific rules. CritiqueVLM: Rule-based scoring criteria are carefully designed, and then we train the CritiqueVLM on human-annotated corpus.

### v3

@@ Abstract
Evaluating the alignment capabilities of large Vision-Language Models (VLMs) is essential for determining their effectiveness as helpful assistants. However, existing benchmarks primarily focus on basic abilities using nonverbal methods, such as yes-no and multiple-choice questions. In this paper, we address this gap by introducing AlignMMBench, which provides more nuanced evaluations of alignment capabilities and is the first benchmark specifically designed for Chinese visual contexts. This benchmark is meticulously curated from real-world scenarios and internet sources, encompassing thirteen specific tasks across three categories, and includes both single-turn and multi-turn dialogue scenarios. Incorporating a prompt rewrite strategy, AlignMMBench encompasses 1,054 images and 4,978 question-answer pairs. To facilitate the evaluation pipeline, we develop CritiqueVLM, a rule-calibrated evaluator that exceeds GPT-4’s evaluation ability. Additionally, we measure the “alignment score”, a quantitative metric designed to assess the robustness and stability of models across diverse prompts. Finally, we evaluate the performance of representative VLMs on AlignMMBench, offering insights into the capabilities and limitations of different VLM architectures. The evaluation code and data are available at https://github.com/THUDM/AlignMMBench.
AlignMMBench: Evaluating Chinese Multimodal Alignment in Large Vision-Language Models
Yuhang Wu1,∗,†, Wenmeng Yu2,∗, Yean Cheng3,†, Yan Wang2, Xiaohan Zhang2, Jiazheng Xu1,†, Ming Ding2, Yuxiao Dong1 1Tsinghua University     2Zhipu AI     3Peking University
Figure 1: Performance vs. Alignment score of various models. Performances are ranged from 0 to 10, while Alignment scores are ranged from 0.2 to $\infty$.
11footnotetext: Equal contributions.22footnotetext: Work was partially done when interned at Zhipu AI.
@@ 1 Introduction
Equipped with Large Language Models (LLMs), Vision-Language Models (VLMs) demonstrate impressive performance in various visual tasks, such as image description and visual question answering Achiam et al. (2023); Wang et al. (2023); Bai et al. (2023a). Following alignment training, like Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), VLMs can comprehend and generate human language within visual contexts.
Recently, significant efforts have been made into developing VLMs for not only English but also Chinese contexts, such as QwenVL Bai et al. (2023a), CogVLM Wang et al. (2023), MiniCPM Hu et al. (2023), Intern-VL Chen et al. (2023b), Intern-XComposer2V Dong et al. (2024), Deepseek-VL Lu et al. (2024a), and Yi-VL AI et al. (2024). Specifically, in publicly available comprehensive rankings [^1] https://rank.opencompass.org.cn/leaderboard-multimodal, these VLMs have demonstrated performance comparable to that of GPT-4o Achiam et al. (2023). To facilitate objective comparisons and evaluations among these models, researchers mainly utilize the following benchmarks: 1) general evaluation sets such as MME Fu et al. (2024) and MMBench Liu et al. (2023a), and 2) domain-specific evaluation sets such as MathVista Lu et al. (2024b) and MMMU Yue et al. (2023). However, these benchmarks primarily assess model capabilities through non-verbal methods and lack detailed evaluations of alignment performance, particularly in the Chinese context.
@@
Ai2D Hiippala et al. (2021) 4,903 English multi-choice Accuracy
LLaVABench Liu et al. (2024) 150 English GPT Score
MathVista Lu et al. (2024b) 6,141 English [^2] MathVista contains 6.57% non-English questions, such as Chinese and Persian. GPT Accuracy
MME Fu et al. (2024) 2,800 English yes-no Accuracy
MMBench Liu et al. (2023a) 3,217 Eng. & Chi. multi-choice Accuracy
MMMU Yue et al. (2023) 11,500 English multi-choice Accuracy
MMStar Chen et al. (2024b) 1,500 English multi-choice Accuracy
MM-Vet Yu et al. (2023b) 205 English GPT Score
OCRBench Liu et al. (2023b) 1,000 English string match Accuracy
TouchStone Bai et al. (2023b) 908 English GPT Score
VisIT-Bench Bitton et al. (2023) 592 English GPT Elo / Win
AlignMMBench (ours) 4,978 Chinese CritiqueVLM Score
@@
Table 1: Comparisons between AlignMMBench and other benchmarks.
@@
However, constructing a high-quality Chinese multimodal alignment benchmark is fully challenging. First, unlike the English context, which benefits from traditional benchmarks such as VQAv2 Goyal et al. (2017), TextVQA Singh et al. (2019), and ChartQA Masry et al. (2022), clean and publicly available Chinese multimodal corpora are exceedingly scarce. It is necessary to collect and construct these resources from scratch. Second, the inherent ambiguity of the Chinese context is more pronounced Huang et al. (1997), necessitating multiple annotators to repeatedly verify a single high-quality Chinese multimodal corpus. Besides, The characteristics of images and the underlying world knowledge can vary significantly between different languages Duncum (2004). As a result, relying solely on English datasets limits the comprehensive evaluation of Chinese vision-language models (VLMs). Therefore, developing a multimodal benchmark within Chinese visual context is essential and promising.
@@
To address these gap, we introduce AlignMMBench, a comprehensive evaluation benchmark specifically designed to assess the Chinese alignment capabilities. Given the scarcity of Chinese multimodal corpora, AlignMMBench is meticulously curated from real-world scenarios and internet resources. It encompasses thirteen specific tasks across three high-level categories, including both single-turn and multi-turn dialogue scenarios, to make detailed evaluation of alignment performance. Considering the diversity of query forms in real-world user scenarios, we introduce a LLM-based prompt-rewritting strategy, which transforms a single query into multiple stylistically distinct yet semantically equivalent questions. Leveraging this strategy, we introduce a new metric, the "alignment score", to investigate the reasons for performance differences among various models. By combining performance metrics with the alignment score, we categorize models into 4 classes, as specified in Figure 1. As shown in Table 1 and Figure 2, AlignMMBench comprises 1,054 images and 4,978 question-answer pairs.
@@
Figure 2: Categories and examples of AlignMMBench. The chart on the left displays the categories of AlignMMBench, encompassing three main categories and thirteen specific tasks. The numbers listed under each category represent the number of images in that category and the corresponding percentage of the total. The right side of the pie chart presents two examples, illustrating instances from the incoherence and coherence tasks.
@@
Due to the lack of standard answers, the evaluation of open-ended questions is a significant and challenging research topic. One simple approach is to utilize the strongest LLM, such as GPT-4 Achiam et al. (2023), to score the model responses. However, the API-based GPT-4 operates as a black box, resulting in limited control for developers. The costs associated with API usage impose limitations on the size of the datasets that can be used as well as the number of models that can be tested. To address these limitations, we fine-tuned an evaluator based on the ChatGLM3-6B Du et al. (2022), named CritiqueVLM. Equipped with the detailed prompt rules and chain-of-thought reasoning, CritiqueVLM achieves a $34.8\%$ reduction in mean absolute error when compared to human scoring, outperforming GPT-4. Additionally, CritiqueVLM features higher consistency across other metrics.
@@
Based on AlignMMBench and CritiqueVLM,we evaluate numerous popular VLMs that support Chinese. Our findings are as follows: (1) All VLMs excel in perception and understanding, achieving an average score of 5.07, but perform poorly in reasoning and analysis, with an average score of 4.38; (2) Within dialogue contexts, VLMs exhibit suboptimal performance in the incoherence task compared to the coherence task, indicating that VLMs struggle to detect previous errors; (3) Certain English-based VLMs, such as Phi-3-Vision Abdin et al. (2024), exhibit suboptimal performance on AlignMMBench, suggesting the composition of the training corpus is critical in alignment evaluation.
The contributions are summarized as follows:
@@
  • We propose AlignMMBench, a multimodal alignment benchmark that encompasses both single-turn and multi-turn dialogue scenarios. It includes three categories and thirteen capability tasks, with a total of 4,978 question-answer pairs. As far as we know, this is the first public alignment benchmark specifically designed for the Chinese visual context.
@@
  • To improve the controllability of evaluation scores, we introduce the CritiqueVLM, a ChatGLM3-6B based evaluator that has been rule-calibrated and carefully fine-tuned. With human judgements, its evaluation consistency surpasses that of GPT-4.
@@
  • We benchmark representative VLMs on AlignMMBench. Beyond their performance, we provide in-depth insights into the current state of Chinese VLM and highlight the areas that require further enhancement.
@@ 2 Related Work
Early multimodal benchmarks predominantly focus on specific cross-modal tasks such as Image Caption Lin et al. (2014); Plummer et al. (2015), Visual Grounding Kazemzadeh et al. (2014); Yu et al. (2016), Visual Question Answering Schwenk et al. (2022); Marino et al. (2019); Mathew et al. (2021) and Optical Character Recognition (OCR) Singh et al. (2021). Recent benchmarks can be categorized into two types based on their questions, including domain-specific and general scenarios. 1) Domain-specific benchmarks concentrate on particular subfields. For example, discipline reasoning Yue et al. (2023); Lu et al. (2024b), OCR Liu et al. (2023b), chart understanding  Masry et al. (2022); Hiippala et al. (2021), and the hallucination test Liu et al. (2024). 2) General benchmarks provide multi-dimensional and diverse assessments, encompassing MME Fu et al. (2024), MMBench Liu et al. (2023a), MMStar Chen et al. (2024b), MMVet Yu et al. (2023b), LLaVABench Chen et al. (2024b), VisIT-Bench Bitton et al. (2023), and TouchStone Bai et al. (2023b). From another perspective, based on differences in result evaluation methods, the aforementioned benchmarks can be divided into deterministic and open-ended types. Table 1 provides an overview of the categories and salient features of prominent benchmarks.
@@
In the realm of automated model response evaluation, employing a robust LLM as an evaluation assistant is widely adopted across various VLM benchmarks. MathVista Lu et al. (2024b) utilizes an LLM to extract answers from responses for more precise comparisons with reference answers, while other works Liu et al. (2024); Yu et al. (2023b); Bai et al. (2023b); Bitton et al. (2023) adopt the LLM-as-a-Judge Zheng et al. (2024) approach. This method prompts the LLM to act as a judge, scoring responses or selecting the better response between two candidates. Furthermore, recent researches Ge et al. (2024); Zhang et al. (2023); Kim et al. (2024, 2023) explore the feasibility of employing the VLM-as-a-Judge approach to achieve more accurate evaluation results.
However, evaluating numerous models within a large benchmark can be costly and inefficient, with results potentially becoming unstable due to API-based model updates. The current capabilities of VLMs in instruction-following and logical reasoning are markedly inferior to those of the most advanced LLMs. In this work, we design a set of task-level evaluation prompts to improve the accuracy of LLM scoring. Additionally, we fine-tune ChatGLM3-6B Du et al. (2022) using a high-quality, human-annotated grading corpus to enhance the evaluation capabilities of our model.
@@ 3.1 Dataset Composition
Figure 3: Overall framework of our work.